OpenStudy (anonymous):

find the fourier series of f(x)=x^2+x+1

OpenStudy (amistre64):

ugh ... ive sat down and watched a video on this for hours and it never sank in lol

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

give me the ved

OpenStudy (anonymous):

have to pass the exam lol

OpenStudy (amistre64):

its an mit lecture .. with a chalkboard :) let me see if i can locate it

OpenStudy (amistre64):

http://freevideolectures.com/Course/2252/The-Fourier-Transform-and-its-Applications fine, its was stanford but we wont fault them for it :)

OpenStudy (anonymous):

ok ok lol

OpenStudy (amistre64):

i found it to be pretty cool, but nothing stuck :/

OpenStudy (anonymous):

really ? well my prof was cool when he explain it xD but i totally forget it

OpenStudy (anonymous):

basically i need to find x then expand for some a_n and try to find pattern

OpenStudy (amistre64):

yeah ... basically lol

OpenStudy (amistre64):

imma gonna have to see if i can download it and make some popcorn ....

OpenStudy (anonymous):

haha lolz xD

OpenStudy (anonymous):

well ty for help :)

OpenStudy (tkhunny):

1) It's neither an even function nor an odd function, so it's not going to be pretty. 2) The Domain of this function is \((-\infty,\infty)\). This is quite a problem for Fourier. Any particular range you would like to repeat in some periodic fashion?

OpenStudy (anonymous):

welli solve it wew ! it took me alote of time xD

OpenStudy (tkhunny):

Really, what periodic assumptions did you make and what was your result?

OpenStudy (anonymous):

x=2 sin x-sin 2x +2/3 sin 3x -1/2 sin 4x .... 1=1 x^2=pi^2/3- 4 cos x + cos 2x - 4/9 cos 3x +1/4 cos 4 x+ ...